6Na + Fe2O3 --> 3Na2O + 2Fe If you are provided 200g of sodium and 250 grams of iron(III) oxide, which substance is the limiting reactant? Explain your reasoning.
1
Expert's answer
2018-03-29T08:00:02-0400
6Na + Fe2O3 → 3Na2O + 2Fe

n(Na) = m/M = 200g/23g/mol = 8.7 moles

n(Fe2O3) = 250g/160g/mol = 1.6 moles

The limiting reactant in a chemical reaction is the substance which is totally consumed when the chemical reaction is complete.

From our calculation, 8.7 moles of Na reacted with only 8.7/6 = 1.45 moles out of 1.6 moles of Fe2O3. Fe2O3 is the excess reactant because what was given is more than what reacted and therefore Na is the limiting reactant.